# Diffusion in Zero-Shot Learning for Environmental Audio

### v1

@@ Abstract
Zero-shot learning enables models to generalize to unseen classes by leveraging semantic information, bridging the gap between training and testing sets with non-overlapping classes. While much research has focused on zero-shot learning in computer vision, the application of these methods to environmental audio remains underexplored, with poor performance in existing studies. Generative methods, which have demonstrated success in computer vision, are notably absent from environmental audio zero-shot learning, where classification-based approaches dominate.
To address this gap, this work investigates generative methods for zero-shot learning in environmental audio. Two successful generative models from computer vision are adapted: a cross-aligned and distribution-aligned variational autoencoder (CADA-VAE) and a leveraging invariant side generative adversarial network (LisGAN). Additionally, a novel diffusion model conditioned on class auxiliary data is introduced. The diffusion model generates synthetic data for unseen classes, which is combined with seen-class data to train a classifier.
Experiments are conducted on two environmental audio datasets, ESC-50 and FSC22. Results show that the diffusion model significantly outperforms all baseline methods, achieving more than 25% higher accuracy on the ESC-50 test partition.
This work establishes the diffusion model as a promising generative approach for zero-shot learning and introduces the first benchmark of generative methods for environmental audio zero-shot learning, providing a foundation for future research in the field.
Code is provided at https://github.com/ysims/ZeroDiffusion for the novel ZeroDiffusion method.
@@ I Introduction
Machine learning has revolutionised how society solves complex problems [1, 2, 3]. Various publications consider the future impact of machine learning, such as AI cities [4], humanoid robotics [5] and medicine [6, 7, 8]. To solve these future problems, machines and AI need to be adaptive [9, 10]. One way machines can be adaptive is to classify new items without labelled training data [11, 12]. This problem is called ‘zero-shot learning’.
@@
Zero-shot learning is a category of supervised machine learning that aims to create a model to classify classes it has not seen during training. That is, the set of classes used in training and testing are disjoint. Zero-shot learning is valuable in real-world applications, such as remote sensing [13], image classification [14, 15], image segmentation [16, 17], environmental audio [18, 19] and security [20], where there may not exist data for every class.
An intuitive way to understand zero-shot learning involves a thought experiment on how humans would solve the problem. Suppose someone has never seen a snake, cat or bird but knows key attributes these animals have. When shown a picture of a cat, the person reasons that it is a cat as it has four legs and whiskers, and no scales or wings. The person performs zero-shot learning using features of the target class and learning what those features mean through knowledge of other animals.
It is important not to confuse zero-shot learning with similar zero-data problems. In zero-shot learning, the method knows about the existence of all classes, i.e. the set of classes is closed. If it was open, i.e. there are unknown classes, the problem is called open-set learning [21, 22, 23]. Domain adaptation [24, 25] is another problem with a lack of data; however, there is no clear ‘seen’ and ‘unseen’ class distinction. An example of domain adaptation is training in simulation and transferring to the real world without fine-tuning [26, 27, 28]. Finally, zero-shot prompting is a newer area that uses large language models to perform new tasks [29, 30, 31]. This differs from zero-shot learning as there is no seen and unseen class distinction, and it is hard to quantify how much overlap the task has to scenarios the model was trained on.
Close to these zero-data problems, particularly zero-shot learning, is few-shot learning. Rather than zero data for select classes, few-shot learning has a small number of samples (e.g. three to five) for particular target classes. A comprehensive review of few-shot learning is provided by Song et al. [32].
In the literature, zero-shot learning methods are commonly tested on computer vision classification tasks [12, 33, 34, 14, 35, 36, 15, 37, 38, 39]. Traditional methods for zero-shot learning in computer vision relied on classification-based models with engineered attributes [12], but more recent methods use generative models [34, 15, 37] such as generative adversarial networks [40] and variational autoencoders [41]. These models have been shown to improve performance on unseen classes and can help mitigate problems such as the hubness problem and the domain shift problem [42]. The hubness problem occurs when data points are concentrated around a minority of classes, leading to poor performance when tested on unseen classes [43]. The domain shift problem arises when the model performs poorly because the training and test sets have different distributions [44].
This work focuses on environmental audio, where zero-shot performance lags considerably behind that of computer vision. It takes inspiration from the successful usage of generative methods in computer vision [14, 45, 34, 37], and from a newer generative method, diffusion, which has seen success in image generation tasks [46, 47]. There is currently no literature known to the authors on environmental audio zero-shot learning using generative methods.
@@
Recent use cases for environmental audio classification include smart cities [48], detection of traffic noise pollution [49], and hearing aid analysis [50]. These use cases motivate research in environmental audio, particularly in the context of limited zero-shot research.
There is a lack of datasets that can be used to benchmark environmental audio zero-shot learning, as datasets are often not released in their raw audio form or lack classes. AudioSet [51] is a valuable large dataset, but instead of audio clips, it contains embeddings. Embeddings are high dimensional vectors representing the data. These embeddings come from a network pre-trained on AudioSet. This pre-training makes the dataset unsuitable for zero-shot learning, as the unseen classes may be present. Other datasets, such as UrbanSound8k [52], have too few classes for rigorous zero-shot learning. Partitioning seen and unseen classes will result in a seen set with little diversity and may not cover a large enough semantic space for the model to generalise to unseen classes, risking overfitting. Additionally, the unseen classes will be very few, and so the statistical significance of the results is poor. A reasonable minimum may be five unseen classes, and double that amount of seen classes, for a total of 15 classes in the dataset.
ESC-50 [53] is an open dataset with fifty classes, often used in environmental audio zero-shot learning literature. FSC22 [54] is another open dataset and contains twenty-seven classes. While this is almost half the size of ESC-50, it is still a substantial size to perform zero-shot learning on. We are not aware of the existence of any zero-shot learning literature using this dataset.
The present study uses ESC-50 and FSC22 to compare three generative methods. These methods use the following architectures: generative adversarial networks (GANs), variational autoencoders (VAEs) and diffusion networks. A bilinear compatibility network [55] is also compared as a baseline. The method ZeroDiffusion is introduced as a new method for zero-shot learning using diffusion.
The GAN method used is LisGAN [15]; however, the component where high-confidence unseen samples are injected back into the training is not included, to prevent violating the zero-shot learning problem formation where unseen class data should not be present during training. The method uses a conditional Wasserstein GAN [56]. Zero-shot learning methods often aim to learn the relationship between feature and class embeddings. To achieve this, the GAN is conditioned on class embeddings. It is then used to generate synthetic samples for unseen classes. Vanilla conditional GAN methods were experimented with in the presented study, but they suffered extreme instability and low performance. The adjusted LisGAN method performs worse than the bilinear compatibility method but is better than a random model.
The cross-aligned and distribution-aligned VAE (CADA-VAE) [14] has been previously used in computer vision with success. The method trains two variational autoencoders, one for auxiliary data and one for feature data, to learn a shared latent space. The results were comparable to the bilinear compatibility method.
Finally, the most successful model is our new diffusion model. The experiments with diffusion show a clear improvement upon any previous method in environmental audio zero-shot learning. While there is existing zero-shot learning literature outside of environmental audio on diffusion models [57, 58], no method trains a network from scratch. The methods use the pre-trained Stable Diffusion model [47]. While this can provide practical usage, it is hard to draw theoretical conclusions on the zero-shot performance of this method. The overlap of the unseen classes and training classes is unknown. This work shows that diffusion, with a transparent training process and a valid seen and unseen class distinction, outperforms other techniques in zero-shot learning.
@@
  • The first diffusion method for zero-shot learning that does not use a pre-trained model.
  • A new state-of-the-art for zero-shot learning with the ESC-50 dataset using the diffusion method.
  • The first benchmark on the FSC22 dataset, with the diffusion method outperforming other methods.
  • The first comparison of GANs, VAEs and diffusion-based methods for environmental audio zero-shot learning with ESC-50 and FSC22.
@@ II Related Work
This section first covers terminology in zero-shot learning to better understand the literature. Then, it examines existing work on generative methods in zero-shot learning, much of which is in the domain of computer vision. Finally, existing works in environmental audio zero-shot learning are considered.
@@ II-A Terminology
In this section, zero-shot learning terminology is presented in an effort to contextualise the literature. First, embeddings for both class and feature data are considered, then the concept of auxiliary data and finally, a high-level definition of zero-shot learning.
@@ II-A1 Embeddings
Embeddings are a method of converting raw data, such as images or audio, into a high-level form that machine learning algorithms can use. The form embeddings take is a high-dimensional vector of real numbers. The methods in this work use deep learning to obtain these vector representations, using both supervised and unsupervised methods. An example of what an embedding space might look like is shown in Figure 1. Lions and cats are both felids and therefore semantically similar and closer to each other in the vector space. ‘Dog’ is semantically closer to the felids than ‘car’ as they are types of animals. In this way, the structure of the space reveals semantic relationships.
@@
Supervised methods to create embeddings typically involve classical supervised learning on a dataset and the output of the second last layer of the artificial neural network becomes the embedded data vector. Zero-shot learning requires a clear distinction between the seen and unseen classes, and so the dataset used to learn the embedding must only contain classes that are not part of the unseen class set. Typically the entire seen portion of the dataset is used to train and test this embedding, as done in the experiments of this paper, but any dataset could be used as long as there is no overlap with the unseen class set.
Unsupervised methods involve learning a mapping to a space where the semantic distance of data is represented in the space. The most common example of this embedding method is word embeddings, such as GloVe [59] and Word2Vec [60]. These methods learn the semantic similarity of words using their proximity in texts. To a human, embeddings may seem to contain little meaning, but these vectors are information-rich. For example, the word embedding vectors for ‘man’, ‘woman’ and ‘king’ can be combined in an equation ‘king’ - ‘man’ + ‘woman’ to approximate the vector for ‘queen’ [60].
@@ II-A2 Auxiliary Data
Zero-shot learning involves learning to classify known classes without samples for those classes. If the classes are unknown, the problem is called open-set learning. These known classes contain information even without samples. This information is called ‘auxiliary data’. The most common form of auxiliary data is the word embedding of the class label, retrieved using Word2Vec or GloVe. Another popular form is human-engineered attributes, where a vector is created for each class and each position in the vector represents whether or not the class has a particular attribute such as ‘has fur’ or ‘eats fish’. The Animals with Attributes [12] and Animals with Attributes 2 [61] datasets are examples of datasets that contain human-engineered attribute vectors, and were created for zero-shot learning research.
@@ II-A3 Zero-Shot Learning
Zero-shot learning involves training on seen classes and testing on unseen classes. Generalised zero-shot learning extends this problem to include both seen and unseen classes in testing. Zero-shot learning is one of many zero-data problems, as described in the introduction.
Feature data is defined as the target data the user wants to classify. In computer vision, the feature data is images. In environmental audio, it is audio clips. Feature data for unseen classes should not be present anywhere before the testing phase, including during pre-training.
The embeddings for the feature data and classes represent semantic information. Zero-shot learning aims to train a model that generalises its understanding of semantics to go beyond class-specific knowledge.
@@ II-B Generative Methods in Zero-Shot Learning
In 2017, autoencoders entered the computer vision zero-shot learning space through two publications, Tsai et al. [33] and Mukherjee et al. [62]. These used two traditional autoencoders for each modality and learnt a shared latent space between these two autoencoders. Conditional variational autoencoders were introduced in zero-shot learning the following year [34]. A new benchmark was found when CADA-VAE [14] was proposed which took advantage of variational autoencoders and created a shared latent space through cross-alignment and distribution-alignment.
@@
One of the first GAN methods in zero-shot learning is f-CLSWGAN, a Wasserstein GAN method [36]. In the same year, another GAN method was created that used a visual pivot regulariser to guide training [63]. It also used Wikipedia text descriptions converted to word embeddings using term frequency-inverse document frequency (TF-IDF). The method was compared with other zero-shot learning methods including a basic GAN model. The basic GAN model performs much worse than any established zero-shot learning method, while Zhu et al.’s method [63] shows an improvement over the state-of-the-art.
@@
LisGAN [15] improved on previous GAN methods by using a Wasserstein GAN conditioned with auxiliary data. They use ‘soul samples’, samples representing the ‘ideal’ sample from a class. Soul samples for seen classes are the average sample over all the seen samples for that class, while soul samples for unseen classes are constructed using the auxiliary data for the class. The GAN is trained to generate data close to a soul sample. The authors perform a step where unseen data is tested and any samples with high confidence are injected into training. This violates the requirement in zero-shot learning where unseen data should not be seen during training, even if it is unlabelled. GDAN [37] proposes a GAN with a regressor and cyclic-consistent loss. The results are marginally better than the conditional VAE method from Mishra et al. [34], indicating that GAN methods still fall behind VAE methods for zero-shot learning, given that at this point better VAE methods had been introduced.
@@
Task Aligned Generative Meta-learning for Zero-shot Learning [38] is a proposed method to reduce bias towards particular classes by aligning the distribution of seen and unseen classes. The backbone of the technique is a meta conditional GAN [64]. SAGAN [39] involves training a GAN, followed by training a classifier to align the visual and semantic space. This method considers what to do with the trained GAN rather than improving the GAN training method. In the same year, ZeroNAS [65] is proposed which performs neural architecture search on GAN zero-shot learning methods to find the optimal architecture for a zero-shot learning problem. The authors claim that GAN methods are hard to train and require experts for each problem scenario, motivating the automation of the expert curation of these networks.
@@
While there is much literature on how GANs may be used for zero-shot learning and how general improvements to GAN training can be used, they are still notoriously hard to train [63, 65]. This does pose the question of if GANs are too unstable and not reliable enough for zero-shot learning problems. In practical situations, it is hard to estimate model performance on unseen classes. Therefore, the method must be reliable.
Chou et al. [35] propose a generative method without VAEs or GANs. It uses a beta distribution to generate virtual classes for more robust zero-shot learning. Despite their efforts, GANs and VAEs still dominate the landscape of zero-shot learning.
@@
In 2023, two publications emerged using diffusion in zero-shot learning [57, 58]. These both exist within the image classification domain and use the pre-trained Stable Diffusion model [47]. These provided insight into the power of diffusion models for zero-shot learning; however, the pre-trained nature of the model and the unknown training set make it hard to measure the ability of diffusion to perform zero-shot learning.

### v2

# Embedding-Space Diffusion for Zero-Shot Environmental Sound Classification

@@ Abstract
Zero-shot learning enables models to generalise to unseen classes by leveraging semantic information, bridging the gap between training and testing sets with non-overlapping classes. While much research has focused on zero-shot learning in computer vision, the application of these methods to environmental audio remains underexplored, with poor performance in existing studies. Generative methods, which have demonstrated success in computer vision, are notably absent from zero-shot environmental sound classification studies.
To address this gap, this work investigates generative methods for zero-shot learning in environmental audio. Two successful generative models from computer vision are adapted: a cross-aligned and distribution-aligned variational autoencoder (CADA-VAE) and a leveraging invariant side generative adversarial network (LisGAN). Additionally, we introduced a novel diffusion model conditioned on class auxiliary data. Synthetic embeddings generated by the diffusion model are combined with seen class embeddings to train a classifier.
Experiments are conducted on five environmental audio datasets, ESC-50, ARCA23K-FSD, FSC22, UrbanSound8k and TAU Urban Acoustics 2019, and one music classification dataset, GTZAN. Results show that the diffusion model outperforms all baseline methods on average across six audio datasets.
This work establishes the diffusion model as a promising approach for zero-shot learning and introduces the first benchmark of generative methods for zero-shot environmental sound classification, providing a foundation for future research.
@@ I Introduction
Machine learning has revolutionised how society solves complex problems [2, 3, 4]. Various publications consider the future impact of machine learning, such as smart cities [5], humanoid robotics [6] and medicine [7, 8, 9]. To solve these future problems, machines and AI need to be adaptive [10, 11]. One way machines can be adaptive is to classify new items without labelled training data [12, 13]. One method to achieve this is zero-shot learning.
@@
Zero-shot learning is a subset of supervised machine learning that aims to create a model to classify classes it has not seen during training. That is, the set of classes used in training and testing are disjoint. Zero-shot learning is valuable in real-world applications, such as remote sensing [14], image classification [15, 16], image segmentation [17, 18], environmental audio [19, 20] and security [21], where there may not exist data for every class.
@@
Outside of zero-shot learning, recent practical use cases for environmental audio classification include smart cities [22], detection of traffic noise pollution [23], and hearing aid analysis [24]. These use cases motivate research in environmental audio, particularly in the context of limited zero-shot research.
An intuitive way to understand zero-shot learning involves a thought experiment on how humans would solve the problem. Suppose someone has never seen a snake, cat or bird but knows key attributes these animals have. When shown a picture of a cat, the person reasons that it is a cat as it has four legs and whiskers, and no scales or wings. The person performs zero-shot learning using attributes of the target class and learning what those attributes mean through knowledge of other animals.
The presented work focuses on environmental audio, where zero-shot performance lags considerably behind that of computer vision. It takes inspiration from the successful usage of generative methods in computer vision [15, 25, 26, 27], and from a newer generative method, diffusion, which has seen success in image generation tasks [28, 29]. There is currently no literature known to the authors on environmental audio zero-shot learning using generative methods.
Six datasets - ESC-50, FSC22, ARCA23K-FSD, TAU Urban Acoustic Scenes 2019, UrbanSound8k and GTZAN - are used to compare four methods - the baseline non-linear compatibility network [30] based on Attribute Label Embedding (ALE) [31], LisGAN [16], CADA-VAE [15] and our new novel method, ZeroDiffusion. This represents a broad study of generative adversarial networks (GANs), variational autoencoders (VAEs) and diffusion for zero-shot learning, with the baseline as a comparison. ZeroDiffusion achieves significant accuracy improvement in most cases.
@@
  • The first diffusion method for zero-shot learning that does not use a pre-trained model, but rather works in the embedding space.
  • A new state-of-the-art method for zero-shot environmental sound classification that performs equally to or outperforms a range of existing methods in zero-shot learning across six audio datasets.
  • The first zero-shot learning benchmark on the FSC22, ARCA23K-FSD and TAU Urban Acoustic Scenes 2019 datasets.
  • The first comparison of GANs, VAEs and diffusion-based methods for zero-shot environmental sound classification with six audio datasets: ESC-50, FSC22, ARCA23K-FSD, UrbanSound8k, TAU Urban Acoustic Scenes 2019 and GTZAN.
@@ II Related Work
This section first covers terminology in zero-shot learning to better understand the literature. Then, it examines existing work on generative methods in zero-shot learning, much of which is in the domain of computer vision. Next, existing works in zero-shot environmental audio classification are considered and audio datasets are covered. Finally, with this context, the novelty and contributions of the presented work are explained.
@@ II-A Terminology
To contextualise the literature, this subsection introduces foundational terminology in zero-shot learning. It first describes class and feature embeddings, followed by the role of auxiliary data. A high-level definition of zero-shot learning is then provided to frame the broader research landscape.
@@ II-A1 Embeddings
Embeddings convert raw data, such as images or audio, into a high-level form that machine learning algorithms can use. The form embeddings take is a high-dimensional vector of real numbers, formally $x\in\mathbb{R}^{\alpha}$ is an example of an embedding with dimension $\alpha$. The methods in this work use deep neural networks to obtain these vector representations, using both supervised and unsupervised methods. An example of what an embedding space might look like is shown in Figure 1. Lions and cats are semantically similar and, therefore, are closer to each other in the vector space than a dog or car. In this way, the structure of the space reveals semantic relationships.
@@
Supervised methods to create embeddings typically involve classical supervised learning on a dataset, and an internal layer of the artificial neural network becomes the embedded data vector. In many cases, the second last layer is used because it is expected to contain high-level meaningful information compared to previous layers.
Zero-shot learning requires a clear distinction between the seen and unseen classes, and so the dataset used to learn the embedding must only contain classes that are not part of the unseen class set. Typically, the entire seen portion of the dataset is used to train and test this embedding, as done in the experiments of this paper, but any dataset is suitable as long as there is no overlap with the unseen class set.
Unsupervised methods involve learning a mapping to a space that represents the semantic distance of data. A common example of this type of embedding method is word embeddings, such as GloVe [32] and Word2Vec [33]. These methods learn the semantic similarity of words using their proximity in texts. To a human, embeddings may seem to contain little meaning, but these vectors are information-rich. For example, the word embedding vectors for ‘man’, ‘woman’ and ‘king’ can be combined in an equation ‘king’ - ‘man’ + ‘woman’ to approximate the vector for ‘queen’ [33]. GloVe uses matrix factorisation and builds a large co-occurrence matrix of word pairs to globally understand the relationship between words, while Word2Vec uses context-based predictive neural networks. While their learning processes are different, both methods have the same goal and provide semantically rich embeddings.
@@ II-A2 Auxiliary Data
Zero-shot learning involves learning to classify known classes without samples for those classes. If the classes are unknown, the problem is open-set learning. These known classes contain information even without samples. This information is ‘auxiliary data’. The most common form of auxiliary data is the word embedding of the class label, retrieved using Word2Vec or GloVe. Another popular form is human-engineered attributes, where a vector is manually created for each class and each position in the vector represents whether or not the class has a particular attribute such as ‘has fur’ or ‘eats fish’. The Animals with Attributes [13] and Animals with Attributes 2 [34] datasets are examples of datasets that contain human-engineered attribute vectors and were created for zero-shot learning research.
@@ II-A3 Zero-Shot Learning
Zero-shot learning involves training a model on a set of seen classes $X$ and evaluating its performance on unseen classes $Y$, where $X\cap Y=\varnothing$.
Feature data refers to the input the model is classifying, such as images in computer vision or audio clips in environmental sound classification. Feature data from unseen classes must not be used before testing, including pre-training.
Both feature data and class labels are represented as embeddings that capture semantic information. The goal of zero-shot learning is to train a model that can leverage these semantic representations to generalise beyond the classes it has been explicitly trained on.
@@ II-B1 Autoencoders
In 2017, autoencoders entered the computer vision zero-shot learning space through two publications, Tsai et al. [35] and Mukherjee et al. [36]. They used two traditional autoencoders for each modality (class textual descriptions and images) and trained a shared latent space between these two modalities. Conditional variational autoencoders were introduced in zero-shot learning the following year [26]. A new benchmark was found when CADA-VAE [15] was proposed, which used variational autoencoders and strengthened the shared latent space through cross-alignment and distribution-alignment of the two modalities.
II-B2 Generative Adversarial Networks
@@
One of the first GAN methods in zero-shot learning is f-CLSWGAN, a Wasserstein GAN method [37]. A Wasserstein GAN is a vanilla GAN with the loss function changed to measure the Wasserstein distance between the real and generated data distributions [38]. It results in more stable training and mitigates mode collapse. In the same year, Zhu et al. introduced another GAN method using a visual pivot regulariser to guide training [39]. The regulariser encourages the generation of features for a class to be close to a fixed pivot point. It used Wikipedia text descriptions converted to word embeddings using the term frequency-inverse document frequency (TF-IDF) as class data. The authors compare their method with other zero-shot learning methods, including a basic GAN model. The basic GAN model has much worse accuracy than any established zero-shot learning method on the vision classification task across benchmark datasets.
@@
LisGAN [16] improved on previous GAN methods by using a Wasserstein GAN conditioned on auxiliary data. They use ‘soul samples’, samples representing the ‘ideal’ sample from a class, to guide generation. The method includes a self-training step where high-confidence samples from the unseen classes are injected into the training process. This makes the approach transductive, since it assumes access to unlabelled unseen data, in contrast to the inductive setting commonly assumed in zero-shot learning where no unseen data is available. GDAN [27] proposes a GAN with a regressor and cyclic-consistent loss. The results are marginally better than the conditional VAE method from Mishra et al. [26]. Given the existence of superior VAE methods, this indicates that GAN methods still fall behind VAE methods for zero-shot learning.
@@
Task Aligned Generative Meta-learning for Zero-shot Learning [40] is a proposed method to reduce bias towards particular classes by aligning the distribution of seen and unseen classes. The backbone of the technique is a meta conditional GAN [41], which combines a meta-learning network and a conditional GAN. SAGAN [42] involves training a GAN, followed by training a classifier to align the visual and semantic space. This method focuses on improving the classifier rather than the GAN. In the same year, ZeroNAS [43] is proposed which performs neural architecture search on GAN-based zero-shot learning methods to find the optimal architecture for a zero-shot learning problem. The authors claim that GAN methods are hard to train and require experts for each problem scenario, motivating the automation of expert curation of these networks.
@@
While there is much literature on using GANs in zero-shot learning, they are still notoriously hard to train [39, 43]. GANs may be too unstable and unreliable for zero-shot learning problems. In practical situations, it is hard to estimate model performance on unseen classes. Therefore, the method must be reliable.
II-B3 Diffusion
@@
In 2023, two publications emerged using diffusion in zero-shot learning [44, 45]. These both exist within the image classification domain and use the pre-trained Stable Diffusion model [29]. These provided insight into the power of diffusion models for zero-shot learning; however, the pre-trained nature of the model and the unknown training set make it hard to measure the ability of diffusion to perform zero-shot learning. Additionally, since Stable Diffusion is foundational to the method, usage is limited to problems in computer vision and the use of textual class auxiliary data.